# Online Feature Updates Improve Online (Generalized) Label Shift Adaptation

### v1

@@ 2. Compute $p_{t+1}=\frac{1}{t}\sum_{\tau=1}^{t}s_{\tau}$
Algorithm 5 Revised UOGD for online feature updates (UOGD-R). See the original version in Equation 9 in (Bai et al., 2022).
@@
Algorithm 6 Revised ATLAS for online feature updates (ATLAS-R). See the original version in Equation 9 in (Bai et al., 2022).
@@ Appendix C The Revision for Previous Online Label Shift Adaptation Algorithms
The revised algorithms to be used in the main algorithm OLS-OFU (Algorithm 1) are FTH-R (Algorithm 4), UOGD-R (Algorithm 5), ROGD-R (Algorithm 3), ATLAS-R (Algorithm 6).
@@ D.1 Theorem for FLHFTL-OFU
Before proving Theorem 2 (in Section 3.2 ) we recall the assumption from Baby et al. (2023) for convenience. We refer the reader to Baby et al. (2023) for justifications and further details of the assumptions.
@@
See 2 Proof:
The algorithm in Baby et al. (2023) requires that the estimate $s_{t}$ in Line 1 of Algorithm 2 is unbiased estimate of the label marginal $q_{t}$. Since $f_{t}^{\prime\prime}$ in Algorithm 2 is independent of the sample $S_{t}$, and since we are working under the standard label shift assumption, due to Lipton et al. (2018) we have that $C_{f_{t}^{\prime\prime},D_{0}^{\prime}}^{-1}\cdot\frac{1}{|S_{t}|}\sum_{x_{t}\in S_{t}}f_{t}^{\prime\prime}(x_{t})$ forms an unbiased estimate of $E_{x\sim\mathcal{P}_{t}^{\text{test}}}[f_{t}^{\prime\prime}(x)]$. Further, from Lipton et al. (2018), the reciprocal of standard deviation of this estimate is bounded below by minimum of the singular values of confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$.
@@
where the second line is due to Jensen’s inequality, third line by Cauchy-Schwartz and last line by Proposition 16 in Baby et al. (2023). This finishes the proof.
@@ D.2 Theorem for ROGD-OFU
We state the assumptions first for the later theorems. These assumptions are similar to Assumption 1-3 in (Wu et al., 2021).
@@
Proof: For any fixed $p$,
@@ D.3 Theorem for FTH-OFU
Proof: Denote $q_{t}:=(\mathcal{P}^{\rm test}_{t}(y=k):k\in[K])$. By the Hoeffding and union bound, we have
@@ D.4 Theorems for UOGD-OFU and ATLAS-OFU
[Regret convergence for UOGD-OFU] Let $f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w)$ denote a network with the same feature extractor as that of $f_{t}^{\prime\prime}$ and a last linear layer with weight $w$. Let $f^{\rm uogd-ofu}=f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t})$, where $w_{t}$ is the weight maintained at round $t$ by Algorithm 5. If we run Algorithm 1 with UOGD in (Bai et al., 2022) and let step size be $\eta$, then under the same assumptions as Lemma 1 in (Bai et al., 2022), UOGD-OFU satisfies that
@@
This guarantee follows from the arguments in Bai et al. (2022) from two basic facts below:
@@
For fact 2, note that the $f_{t}^{\prime\prime}$ only uses the data until round $t-1$. So by the same arguments in Bai et al. (2022), using the BBSE estimator defined from the classifier $f_{t}^{\prime\prime}$, the unbiased estimate of risk gradient can be defined.
@@
Rest of the proof is identical to Bai et al. (2022).
@@
Let $f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w)$ denote a network with the same feature extractor as that of $f_{t}^{\prime\prime}$ and a last linear layer with weight $w$. Let $f^{\rm atlas-ofu}=f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t})$, where $w_{t}$ is the weight maintained at round $t$ by Algorithm 6. If we run Algorithm 1 with ATLAS in (Bai et al., 2022) and set up the step size pool =$\mathcal{H}=\{\eta_{i}=O\left(\frac{\sigma}{\sqrt{KT}}\right)\cdot 2^{i-1}|i\in[N]\}$ ($N=1+\lceil\frac{1}{2}\log_{2}(1+2T)\rceil$), then under the same assumptions as Lemma 1 in (Bai et al., 2022), UOGD-OFU satisfies that
@@
The proof is similar to that of Theorem 5 and hence omitted.
@@ Appendix E Additional experiments
E.1 Additional details of datasets
@@
For each type of corruption in CIFAR-10C, we select a mild level and a high level of severity in the experiment section. Here we introduce the exact parameters of mild and high levels of severity for those corruptions. For Gaussian Noise, the severity levels for [mild, high] are [0.03, 0.07]. For Fog, the severity levels for [mild, high] are [(0.75,2.5), (1.5,1.75)]. For Pixelate, the severity levels for [mild, high] are [0.75, 0.65].
Details of additional datasets.
In addition to CIFAR-10 and CIFAR-10C, we evaluate on three more datasets: STL10 (Coates et al., 2011), CINIC (Darlow et al., 2018), and EuroSAT (Helber et al., 2019). Similar to CIFAR-10, we split the original train sets of these datasets into the train set and the validation set by the ratio $4:1$ and use the original test sets for sampling test images in the online test stage.
E.2 Results on Additional Datasets
Figure 2(b) and Figure 5 show the results for three additional datasets: STL10, CINIC, and EuroSAT, also with OLS-OFU (rotation degree prediction) under sinusoidal shift setting. In-line with CIFAR-10, we observe that OLS and OLS-OFU can perform better than Base and OFU, and OLS-OFU can outperform OLS. We find this pattern to be more consistent for EuroSAT and STL10 than CINIC.
Figure 5: Results on three more datasets. SSL in OLS-OFU is rotation degree prediction and online shift pattern is Bernoulli shift.
E.3 More Results on CIFAR-10
Figure 6: Results of Bernoulli shifts on CIFAR-10. OLS-OFU is evaluated with three SSL methods.
Figure 6 shows the results on CIFAR-10 for Bernoulli shift cross three SSL methods in OLS-OFU. Similar to Figure 2(a), OLS-OFU within all three SSL methods outperforms all baseline methods.
E.4 More Results on CIFAR-10C
We evaluate three SSL methods in OLS-OFU on CIFAR-10C for two online shift patterns. We pick moderately high severity levels for evaluating CIFAR-10C. In Figure 7 we can observe the consistent improvement from OLS to OLS-OFU and the SOTA performance of OLS-OFU.
E.5 More Results on CIFAR-10C with High Severity
Figure 8: Results of two online shift patterns on CIFAR-10C (high severity) and three SSL methods in OLS-OFU.
We evaluate three SSL methods in OLS-OFU on CIFAR-10C with high severity for two online shift patterns. In Figure 8, we can observe when SSL in OLS-OFU is rotation degree prediction or MoCo, the improvement from OLS to OLS-OFU is very significant but OLS-OFU cannot outperform OFU. The conclusion is similar to the discussion for Figure 4 in Section 4. However, OLS-OFU with SSL entropy minimization has different behavior. When the corruption of CIFAR-10C becomes more severe, OLS-OFU entropy minimization shows less improvement from OLS, if we compare the transition: Figure 7 (clean CIFAR-10), Figure 7 (CIFAR-10C with mild severity), Figure 8 (CIFAR-10C with high severity). This suggests that rotation degree prediction and MoCo are more appropriate SSL to address the domain shift from CIFAR-10 to CIFAR-10C.
E.6 Empirical evaluation for Equation 8
Figure 9 shows the comparison between LHS (OLS-OFU) and RHS (OLS) of the inequality in Equation 8 when the SSL in OLS-OFU is entropy minimization or MoCo. Similar to what we observe in Figure 3, the inequality in Equation 8 holds cross 4 data settings and two online shift patterns when OLS-OFU is implemented with entropy minimization or MoCo.
Figure 9: Sanity check for the inequality $\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;{\color[rgb]{0.84,0.09,0.41}f_{t}^{\prime\prime}},q_{t}/q_{0});\mathcal{P}_{t}^{\rm test})\right]<\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{0},q_{t}/q_{0});\mathcal{P}_{t}^{\rm test})$ (Equation 8) when SSL in OLS-OFU is entropy minimization or MoCo.
E.7 Ablation for the order of updates and predictions
In the default framework of online distribution shift as shown in Figure 1, the model updates happen after the predictions for the samples at time step $t$. We would like to see if the model updated before the predictions would bring benefit. Figure 10 shows the comparison between ”predict first” and ”update first”. We can observe that there is no strong evidence to demonstrate the advantage of ”predict first” or ”update first” – the difference is indeed insignificant. However, because in the framework of ”predict first” OLS and OLS-OFU enjoy strong theoretical guarantees, we recommend ”predict first” in practice.
Figure 10: Ablation for the order of updates and predictions.
E.8 Details of SSL Methods
@@
When the SSL loss is rotation degree prediction, it requires another network $f^{\rm deg}$ to predict the rotation degree, sharing the same feature extractor $\theta^{\rm feat}$ as $f_{0}$ but with a different set of downstream layers. Its SSL loss $\ell_{\rm ssl}(S;f)$ is defined as $\sum_{x\in S}\ell_{ce}(f^{\rm deg}(R(x,i)),i)$, where $i$ is an integer uniformly sampled from $[4]$, and $R(x,i)$ is to rotate $x$ with degree $\mathrm{DL}[i]$ from a list of degrees $\mathrm{DL}=[0,90,180,270]$. Alternatively, if the SSL loss is entropy minimization, $\ell_{\rm ssl}(S;f)$ would be the entropy $\sum_{x\in S}\sum_{k=1}^{K}f(x)_{k}\log f(x)_{k}$. Moreover, the SSL loss would be a contrastive loss (InfoNCE) where the positive example $x^{\prime}$ is an augmented version of $x$ and other samples in the same time step can be the negative examples. However, the batch size for a time step is small, e.g. $10$ in our experiment. MoCo updates with such a small batch won’t work. Thus, we experiment with MoCo by applying a batch accumulation strategy, which we introduce next.
E.8.1 Batch accumulation for MoCo
Prior test-time training (rotation degree prediction, entropy minimization) methods operate with the OFU framework easily, with feature extractor updates taking place in each time step. Rotation degree prediction originates as a self-supervised training method (Gidaris et al., 2018), thus naturally we evaluate more recent self-supervised training methods, namely MoCo (He et al., 2020; Chen et al., 2020b, 2021). No prior work shows how to use MoCo (or self-supervised learning in general) in a test-time training setting. Given that self-supervised training is sensitive to batch size, the intuition is that a larger batch size (much larger than the number of online samples available per time step) is required to perform a valid gradient update for a MoCo checkpoint. As such, we evaluated a batch accumulation strategy OLS-OFU (BA=$\tau$), where we continue evaluating online samples per time step, but only perform the online feature update at every $\tau$ steps (having accumulated the online samples throughout the $\tau$ steps in one batch). In particular, we perform feature extractor update every $\tau=50$ steps (for 1000 steps, feature update occurs 20 times, online samples evaluation occurs 1000 times), evaluating with 10 online samples per time step, using a smaller learning rate (0.0005) but test-time train with 10 epochs. Notice that $\tau=1$ is the default setting in Algorithm 1. OLS-OFU with MoCo presented in Figure 2, Figure 6 and Figure 7 is equivalent to OLS-OFU (BA=$50$).
To show the necessity of large $\tau$, we evaluate both OLS-OFU (BA=1) and OLS-OFU (BA=50) on CIFAR-10 (Figure 11) and CIFAR-10C (Figure 12). Firstly, we can observe that OLS-OFU (BA=1) is even worse than OLS. We hypothesize this is because small batch size of MoCo will hurt the performance and larger batch size in MoCo is necessary. Hence, we increase $\tau$ from $1$ to $50$ and then we can observe the significant improvement from OLS-OFU (BA=1) to OLS-OFU (BA=50). Now, OLS-OFU (BA=50) can outperform OLS.
Figure 11: Evaluating OLS-OFU with different batch accummulations for MoCo on CIFAR-10.
Figure 12: Evaluating OLS-OFU with different batch accummulations for MoCo on CIFAR-10C.
E.9 Improvements compared with literature
We would like to compare the improvements in the literature for the OLS problem and the improvements from the SOTA in the literature to our method OLS-OFU. We can calculate the difference $\Delta_{\rm error}$ between the errors of method A and B to show the improvement $A\to B$. To measure the improvements in the literature for the OLS problem, we measure $\Delta_{\rm error}$ between FTH (proposed in the first OLS paper) and FLHFTL (SOTA in the literature). We also calculate $\Delta_{\rm error}$ between FLHFTL and FLHFTL-OFU.
[…]
Figure 13: Improvements from FTH to FLHFTL and improvements from FLHFTL to FLHFTL-OFU across all datasets and online shift patterns.
E.10 Self-training

### v3

@@ 2. Compute $p_{t+1}=\frac{1}{t}\sum_{\tau=1}^{t}s_{\tau}$
Algorithm 4 Revised UOGD for online feature updates (UOGD-R). See the original version in Equation 9 in [7].
@@
Algorithm 5 Revised ATLAS for online feature updates (ATLAS-R). See the original version in Equation 9 in [7].
@@
     Output at time $t$: $f_{t+1}^{\prime}$.
  end for
Algorithm 6 Revised FLHFTL for online feature updates (FLHFTL-R); See the original version in [6].
  Require: Online regression oracle ALG.
  for $t=1,\cdots,T$ do
[…]
     2. Compute $\tilde{q}_{t+1}:={\rm ALG}(s_{1},\cdots,s_{t})$
[…]
@@ Appendix C The Revision for Previous Online Label Shift Adaptation Algorithms
The revised algorithms to be used in the main algorithm OLS-OFU (Algorithm 1) are FTH-R (Algorithm 3), UOGD-R (Algorithm 4), ROGD-R (Algorithm 2), ATLAS-R (Algorithm 5), FLHFTL-R (Algorithm 6).
@@ D.1 Theorem for FLHFTL-OFU
Before proving Theorem 2 (in Section 3.3 ) we recall the assumption from Baby et al. [6] for convenience. We refer the reader to Baby et al. [6] for justifications and further details of the assumptions.
@@
Theorem 2.
[Regret convergence for FLHFTL-OFU] Suppose we choose the OLS-R to be FLHFTL-R (Algorithm 6) from Baby et al. [6]. Let $f_{t}^{\rm flhftl-ofu}$ be the output at time step $t-1$ from Algorithm 1, that is $g(\cdot;f_{t}^{\prime\prime},\frac{\tilde{q}_{t}}{q_{0}})$. Let $\sigma$ be the smallest among the the minimum singular values of invertible confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$. Then under Assumptions 1 and 2 in Baby et al. [6], FLHFTL-OFU has the guarantee below:
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm flhftl-ofu};\mathcal{P}_{t}^{\rm test})\right]\leq\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})\right]+O\left(\frac{K^{1/6}V_{T}^{1/3}}{\sigma^{2/3}T^{1/3}}+\frac{K}{\sigma\sqrt{T}}\right),$ (8)
Proof: The proof follows the similar idea in the original online regression algorithm FLHFTL [22] and its variants for solving online label shift problem in Baby et al. [6].
The algorithm in Baby et al. [6] requires that the estimate $s_{t}$ in Line 1 of Algorithm 6 is unbiased estimate of the label marginal $q_{t}$. Since $f_{t}^{\prime\prime}$ in Algorithm 6 is independent of the sample $S_{t}$, and since we are working under the standard label shift assumption, due to Lipton et al. [35] we have that $C_{f_{t}^{\prime\prime},D_{0}^{\prime}}^{-1}\cdot\frac{1}{|S_{t}|}\sum_{x_{t}\in S_{t}}f_{t}^{\prime\prime}(x_{t})$ forms an unbiased estimate of $E_{x\sim\mathcal{P}_{t}^{\text{test}}}[f_{t}^{\prime\prime}(x)]$. Further, from Lipton et al. [35], the reciprocal of standard deviation of this estimate is bounded below by minimum of the singular values of confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$.
@@
where the second line is due to Jensen’s inequality, third line by Cauchy-Schwartz and last line by Proposition 16 in Baby et al. [6]. This finishes the proof.
@@ D.2 Theorem for ROGD-OFU
We state the assumptions first for the later theorems. These assumptions are similar to Assumption 1-3 in [49].
@@
Proof: The proof follows the similar idea in the original online leanring algorithm online gradient descent (OGD) [44] and the ROGD in Wu et al. [49]. For any fixed $p$,
@@ D.3 Theorem for FTH-OFU
Proof: The proof follows the similar idea in the original online leanring algorithm FTL [44] and the FTH in Wu et al. [49]. Denote $q_{t}:=(\mathcal{P}^{\rm test}_{t}(y=k):k\in[K])$. By the Hoeffding and union bound, we have
@@ D.4 Theorems for UOGD-OFU and ATLAS-OFU
[Regret convergence for UOGD-OFU] Let $f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w)$ denote a network with the same feature extractor as that of $f_{t}^{\prime\prime}$ and a last linear layer with weight $w$. Let $f^{\rm uogd-ofu}=f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t})$, where $w_{t}$ is the weight maintained at round $t$ by Algorithm 4. If we run Algorithm 1 with UOGD in [7] and let step size be $\eta$, then under the same assumptions as Lemma 1 in [7], UOGD-OFU satisfies that
@@
This guarantee follows from the arguments in Bai et al. [7] from two basic facts below:
@@
For fact 2, note that the $f_{t}^{\prime\prime}$ only uses the data until round $t-1$. So by the same arguments in Bai et al. [7], using the BBSE estimator defined from the classifier $f_{t}^{\prime\prime}$, the unbiased estimate of risk gradient can be defined.
@@
Rest of the proof is identical to Bai et al. [7].
Theorem 5 for UOGD-OLS shows a bound that depends on the learning rate and one can set up the learning carefully to get the optimal rate. However, this learning rate requires the prior information of $V_{T}$, which is typically unknown during the learning process. ATLAS-OLS overcomes this issue by applying the same online ensembling framework [54] as the original ATLAS. Without knowing this prior information, ATLAS-OLS is able to achieve the same rate of upper bound too, which is stated in the following theorem.
@@
Let $f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w)$ denote a network with the same feature extractor as that of $f_{t}^{\prime\prime}$ and a last linear layer with weight $w$. Let $f^{\rm atlas-ofu}=f(\cdot;\theta^{\rm feat}_{f_{t}^{\prime\prime}},w_{t})$, where $w_{t}$ is the weight maintained at round $t$ by Algorithm 5. If we run Algorithm 1 with ATLAS in [7] and set up the step size pool =$\mathcal{H}=\{\eta_{i}=O\left(\frac{\sigma}{\sqrt{KT}}\right)\cdot 2^{i-1}|i\in[N]\}$ ($N=1+\lceil\frac{1}{2}\log_{2}(1+2T)\rceil$), then under the same assumptions as Lemma 1 in [7], ATLAS-OFU satisfies that
@@
The proof is similar to that of Theorem 5, which mainly follows the similar idea in the original online ensembling framework [54] and ATLAS [7], and hence omitted.
@@ Appendix E Additional experiments
Figure 4: Evaluation of OLS and OLS-OFU.
E.1 Additional Details of Datasets, Online Shift Patterns, and SSL
@@
For each type of corruption in CIFAR-10C, we select a mild level of severity in the experiment section. Here we introduce the exact parameters of mild and high levels of severity for those corruptions. For Gaussian Noise, the severity level is 0.03. For Fog, the severity level is (0.75,2.5). For Pixelate, the severity level is 0.75.
Details of online shift patterns.
Specifically, given two label distribution vectors $q$ and $q^{\prime}$, we simulate the label marginal distributions at time $t$ as a weighted combination of them: $q_{t}:=\alpha_{t}q+(1-\alpha_{t})q^{\prime}$. In Sinusoidal shift, $\alpha_{t}=\sin\frac{i\pi}{L}$ (periodic length $L=\sqrt{T}$, $i=t\textnormal{ mod }L$) while in Bernoulli shift, $\alpha_{t}$ is a random bit (either 0 or 1), where the bit switches $\alpha_{t}=1-\alpha_{t-1}$ with probability $p=1-\frac{1}{\sqrt{T}}$. In our experiments, we set the initial distribution vector $q$ and $q^{\prime}$ with $\frac{1}{K}(1,\cdots,1)$ and $(1,0,\cdots,0)$. To sample the batch test data at time $t$, we first sample a batch of labels (not revealed to the learner) according to $q_{t}$. Then given each label we can sample an image from the test set, and collect this batch of images without labels as $S_{t}$.
Details of self supervised learning techniques.
rotation degree prediction involves initially rotating a given image by a specific degree from the set $\{0,90,180,270\}$ and the classifier is required to determine the degree by which the image has been rotated. Entropy minimization utilizes a minimum entropy regularizer, with the motivation that unlabeled examples are mostly beneficial when classes have a small overlap. MoCo is a more advanced representation learning technique, using a query and momentum encoder to learn representations from unlabeled data by maximizing the similarity of positive pairs and minimizing the similarity of negative pairs.
@@
Their loss functions are as follow. When the SSL loss is rotation degree prediction, it requires another network $f^{\rm deg}$ to predict the rotation degree, sharing the same feature extractor $\theta^{\rm feat}$ as $f_{0}$ but with a different set of downstream layers. Its SSL loss $\ell_{\rm ssl}(S;f)$ is defined as $\sum_{x\in S}\ell_{ce}(f^{\rm deg}(R(x,i)),i)$, where $i$ is an integer uniformly sampled from $[4]$, and $R(x,i)$ is to rotate $x$ with degree $\mathrm{DL}[i]$ from a list of degrees $\mathrm{DL}=[0,90,180,270]$. Alternatively, if the SSL loss is entropy minimization, $\ell_{\rm ssl}(S;f)$ would be the entropy $\sum_{x\in S}\sum_{k=1}^{K}f(x)_{k}\log f(x)_{k}$. Moreover, the SSL loss of MoCo would be a contrastive loss (InfoNCE) where the positive example $x^{\prime}$ is an augmented version of $x$ and other samples in the same time step can be the negative examples.
E.2 Additional Results of OLS-OFU and Baselines
Figure 2 has three subfigures, showing the results of OLS-OFU with three SSL on CIFAR-10, the results on three more datasets for evaluating online label shift, and the results on CIFAR-10C for evaluating online generalized label shift; the online shift patterns are Sinusoidal shift for all these figures in Figure 2. We show similar figures in Figure 4 but now the online shift patterns are Bernoulli shift. We can have the exact same observations from Figure 4: the improvements from original OLS methods to OLS-OFU are significant and consistent across 6 OLS methods, 3 SSL techniques, 4 datasets with online label shift and 3 datasets with online generalized label shift.
We also evaluate three SSL methods in OLS-OFU on CIFAR-10C with mild severity for two online shift patterns. In Figure 7, the improvement from OLS to OLS-OFU is very significant but OLS-OFU cannot outperform OFU.
E.3 Ablation Study on Frequency $\tau$ in OLS-OFU for Entropy Minimization and MoCo
Table 1 in the main paper shows the ablation study on $\tau$ when SSL is rotation degree prediction. Here we include Table 4 and Table 5 for the same ablation study on $\tau$ but the SSL in OLF-OFU is Entropy Minimization and MoCoin two tables respectively. We have same observations as what we observed from Table 1: 1. OLS-OFU with all $\tau$ outperforms OLS; 2. The average error of OLS-OFU decreases from $\tau=1$ (w/o batch accumulation) to $\tau=100$ and starts to increase after and this is because the effectiveness of SSL benefits from $\tau>1$ while larger $\tau$ means less updates and hence hurts the long term performance; 3. the additional time cost of OLS-OFU with larger $\tau$ is smaller and $\tau=100$ gives a great balance of performance and time cost.
FTFWH FTH ROGD ATLAS UOGD FLH-FTL
OLS 11.52% / 0.397 11.56% / 0.395 14.21% / 1.15 12.02% / 42.5 11.65% / 312 11.63% / 1.09
OLS-OFU-BA ($\tau=1$) 10.5% / 260 10.5% / 264 14./ 380 12.1% / 44.2 11.6% / 321 10.6% / 287
OLS-OFU-BA ($\tau=10$) 9.19% / 28.1 9.64% / 28.7 13.2% / 40.9 12% / 43.9 10.6% / 319 9.39% / 28.4
OLS-OFU-BA ($\tau=50$) 7.54% / 6.16 8.41% / 6.3 12.1% / 9.26 11.7% / 43.8 9.4% / 317 7.82% / 6.88
OLS-OFU-BA ($\tau=100$) 6.79% / 3.4 7.57% / 3.38 10.9% / 5.21 10.5% / 43.6 8.46% / 314 7.04% / 4.09
OLS-OFU-BA ($\tau=500$) 11.2% / 1.19 11.3% / 1.18 14.1% / 2.22 11.9% / 42.7 11.4% / 314 11.3% / 1.88
Table 4: Average error / time (minutes) of 6 original OLS methods versus OLS-OFU with various frequency $\tau$ in batch accumulation. The SSL in OLS-OFU is Entropy Minimization.
FTFWH FTH ROGD ATLAS UOGD FLH-FTL
OLS 6.21% / 0.392 7.03% / 1.05 8.23% / 0.888 7.82% / 42.2 6.97% / 302 6.45% / 0.44
OLS-OFU-BA ($\tau=1$) 6.76% / 277 8.13% / 280 8.86% / 374 7.57% / 56.3 7.84% / 319 8.55% / 271
OLS-OFU-BA ($\tau=10$) 5.64% / 28.6 7.08% / 29.7 8.01% / 39.7 7.2% / 44.7 6.87% / 313 7.2% / 28.5
OLS-OFU-BA ($\tau=50$) 4.52% / 6.18 6.03% / 6.96 7.16% / 8.77 6.84% / 43.7 5.9% / 312 5.86% / 6.09
OLS-OFU-BA ($\tau=100$) 3.97% / 3.42 5.3% / 4.2 6.3% / 4.91 6.01% / 43.4 5.19% / 308 5.15% / 3.36
OLS-OFU-BA ($\tau=500$) 5.06% / 2.13 6.32% / 2.86 7.52% / 3.2 7.12% / 43.1 6.23% / 309 6.05% / 2.12
Table 5: Average error / time (minutes) of 6 original OLS methods versus OLS-OFU with various frequency $\tau$ in batch accumulation. The SSL in OLS-OFU is MoCo.
E.4 Empirical Validation of Equation 7
Similar to Figure 3, as shown in Figure 5, it is evident that OLS-OFU with SSL chosen as Entropy Minimization and MoCo is able to yield improvements on the baseline of the regret as shown in Equation 7.
E.5 Self-training
Figure 6: Results on pseudo-labelling.